### v1

# HNN extensions of free groups with equal associated subgroups of finite index: polynomial time word problem

Abstract.
Let $G=F\ast_{\varphi}t$ be an HNN extension of a free group $F$ with two equal associated normal subgroups $H_{1}=H_{2}$ of finite index. We prove that the word problem in $G$ is decidable in polynomial time. This result extends to the case where the subgroups $H_{1}=H_{2}$ are not normal, provided that the isomorphism $\varphi:H_{1}\to H_{2}$ satisfies an additional condition described in Section 5.
Keywords. HNN extensions of free groups, word problem, complexity.
2020 Mathematics Subject Classification. 20F10, 68W30.
1. Introduction
The study of computational problems in the theory of groups began in the early twentieth century. Two central themes in this area are decidability and computational complexity, that together shape our understanding of which problems can be solved algorithmically and how efficiently. In his 1911 work [4], M. Dehn introduced three fundamental decision problems: the word problem, the conjugacy problem, and the isomorphism problem, that have since been central to the field. A significant result concerning decidability was established in the 1960s when Novikov [13] and Boone [1] demonstrated the existence of finitely presented groups for which the word problems are undecidable. Nevertheless, for many important classes of groups, such as automatic groups, finitely generated linear groups, and finitely presented residually free groups, the word problem remains decidable.
The 1940s marked the introduction of HNN extensions by G. Higman, B. Neumann, and H. Neumann [6], providing a powerful tool for group embeddings and for constructing groups with special algorithmic properties, where the word problem is typically decidable. Subsequent research in the 1970s, notably by C. Miller et al [12], further explored the computational complexity of HNN extensions of free groups. This led to the construction of Miller’s machine, a group exhibiting a decidable word problem but an undecidable conjugacy problem.
1.1. HNN extensions
Let $G={\left\langle X\mid R\right\rangle}$, $H_{1},H_{2}\leq G$ and $\varphi:H_{1}\rightarrow H_{2}$ be a group isomorphism. The HNN extension of $G$ relative to $\varphi$ is the group denoted by $G\ast_{\varphi}t$, given by the following presentation:
$G\ast_{\varphi}t={\left\langle X,t\mid R,~t^{-1}ht=\varphi(h),~h\in H_{1}\right\rangle}.$
It is easy to see that if $H_{1}={\left\langle h_{1},\ldots,h_{k}\right\rangle}$, then
$G\ast_{\varphi}t={\left\langle X,t\mid R,~t^{-1}h_{1}t=\varphi(h_{1}),\ldots,t^{-1}h_{k}t=\varphi(h_{k})\right\rangle}.$
For the group $G\ast_{\varphi}t$
  • the group $G$ is called the base group,
  • $t$ is called the stable letter,
  • $H_{1}$ and $H_{2}$ are called the associated subgroups.
Elements of $G\ast_{\varphi}t$ can be defined as alternating sequences of the form
(1) $w=w_{0}t^{\varepsilon_{1}}w_{1}\dots w_{k-1}t^{\varepsilon_{k}}w_{k},$
where $w_{0},\dots,w_{k}$ are group-words over the alphabet of $G$, called syllables, and $\varepsilon_{i}=\pm 1$. The number $k$ is called the syllable length of $w$.
We say that $w$ is $t$-reduced if it is reduced and does not involve the following subwords:
  • $t^{-1}w_{i}t$, where $w_{i}\in H_{1}$;
  • $tw_{i}t^{-1}$, where $w_{i}\in H_{2}$.
Otherwise, we say that $w$ is not $t$-reduced. If $w$ is not $t$-reduced, then it can be simplified as follows:
  • $t^{-1}w_{i}t$, where $w_{i}\in H_{1}$, can be replaced with $\varphi(w_{i})$;
  • $tw_{i}t^{-1}$, where $w_{i}\in H_{2}$, can be replaced with $\varphi^{-1}(w_{i})$.
These operations are called $t$-reductions (or Britton reductions). They do not change the corresponding group element and decrease the syllable length of $w$. Hence, in finitely many steps one obtains an equivalent $t$-reduced word.
Lemma 1.1 (Britton’s lemma, [3]).
$w=1$ in $G\ast_{\varphi}t$ and $k\geq 1$ $\Rightarrow\ \ w$ is not $t$-reduced.
Corollary 1.2.
If the membership problem for $H_{1}$ and $H_{2}$ is decidable, $\varphi$ and $\varphi^{-1}$ are computable, and the word problem for $G$ is decidable, then the word problem for $G\ast_{\varphi}t$ is decidable.
Current state of knowledge regarding the computational properties of the word problem for HNN extensions of free groups can be summarized as follows.
  • The word problem, when approached via Britton’s lemma [3], has exponential-time complexity.
  • In the generic (typical) case, the conjugacy problem can be solved in polynomial time [2, 16].
  • For ascending HNN extensions (when one of the subgroups is the entire group $G$) Lohrey [10] established polynomial-time decidability using straight-line programs.
  • N. Haubold and M. Lohrey [5] also proved that the compressed word problem for an HNN-extension with A finite is polynomial time Turing-reducible to the compressed word problem for the base group H.
  • A special case with equal subgroups associated by the identity isomorphism can be solved in polynomial time [15].
The main computational challenge of Britton reduction is that a single reduction step can multiply the length of a word by a constant factor, potentially producing words of exponential length. We address this issue by representing such exponentially long words using straight-line programs (reviewed in Section 3) that define paths in the subgroup graphs of $H_{1}$ and $H_{2}$ (reviewed in Section 2).
1.2. Our results
The main contributions of this paper are summarized in the following theorems.
Theorem 4.3.
Suppose that $H_{1}=H_{2}$ are normal subgroups of $F$ of finite index and let $\varphi:H_{1}\to H_{2}$ be an isomorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
Theorem 4.3 can be generalized to the case where $H_{1}=H_{2}$ are subgroups of $F$ of finite index and $\varphi$ can be restricted to an isomorphism $\varphi:N\to N$ of a normal subgroup $N\unlhd F$ of finite index. We call such $\varphi$ normalizable in Section 5.
Theorem 5.15.
Suppose that $H_{1}=H_{2}$ are subgroups of $F$ of finite index and let $\varphi:H_{1}\to H_{2}$ be a normalizable isomorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
1.3. Outline
The paper is organized as follows. Section 2.1 introduces essential preliminaries of free groups and subgroup graphs. In Section 3 we discuss the definition and basic properties of straight-line programs. Section 4 presents a polynomial-time algorithm for the word problem in $F\ast_{\varphi}t$ in the case where $H_{1}=H_{2}$ are normal subgroups of finite index, which establishes Theorem 4.3. Section 5 defines the property of $\varphi:H_{1}\to H_{2}$ to be normalizable, presents a polynomial-time algorithm for the word problem in $F\ast_{\varphi}t$ in the case where $H_{1}=H_{2}$ are subgroups of finite index and $\varphi$ is normalizable, which establishes Theorem 5.15.
1.4. Model of computation and internal data representation
We assume that all computations are performed on a random access machine. Data representation for words is discussed in Section 2.1.1 and data representation for straight-line programs is discussed in Section 3.2.
2. Preliminaries: subgroup graphs
2.1. Free groups and free monoids
Recall that an alphabet $X=\{x_{1},\dots,x_{n}\}$ is a set, whose elements are called symbols. For $x\in X$ define the symbol $x^{-1}$ called the inverse of $x$, define the set $X^{-}=\left\{\,x^{-1}\;\middle|\;x\in X\,\right\}$, and form a symmetrized alphabet (group alphabet) $X^{\pm}=X\cup X^{-}$. We refer to elements of $X$ as positive letters and elements of $X^{-}$ as negative letters. The operation -1 defines an involution on the set $X^{\pm}$, mapping each $x\in X$ to $x^{-1}\in X^{-1}$ and $x^{-1}\in X^{-1}$ back to $x\in X$.
A word over the alphabet $X$ is a sequence of letters from $X$. The empty sequence of letters (the empty word) is denoted by $\varepsilon$. In our notation for words, we omit commas between letters and simply write $w=x_{1}\dots x_{n}$. The set of all words over the alphabet $X$ is denoted by $X^{\ast}$. The set $X^{\ast}$ equipped with the binary operation of concatenation is a free monoid.
A group word $w$ is a word over a group alphabet $X^{\pm}$. We use the following notation for group words:
$w=x_{1}^{\varepsilon_{1}}\dots x_{n}^{\varepsilon_{n}}$
where $x_{i}\in X$ and $\varepsilon_{i}=\pm 1$. We say that $w$ is reduced if it does not contain any pair of consecutive inverse letters, that is, any subword of the form $xx^{-1}$ or $x^{-1}x$. Denote by $F(X)$ the set of all reduced group words over $X$. Every word $w$ can be reduced by a process called reduction which successively removes occurrences of subwords of the form $xx^{-1}$ or $x^{-1}x$ until no such subwords remain. The result of reducing any word $w$ is uniquely defined, that is, it does not depend on a particular sequence of removals. Denote by $\overline{w}$} the result of reducing $w$. The set $F(X)$ equipped with the multiplication operation $\cdot$ defined by
$u\cdot v=\overline{u\circ v}$
is a free group. In this paper we mainly consider group words, and for simplicity, we refer to them as words.
2.1.1. Data representation for words
A positive letter $x_{i}$ of an alphabet $X=\{x_{1},\dots,x_{n}\}$ is encoded by $i\in{\mathbb{Z}}$ and a negative letter $x_{i}^{-1}$ is encoded by $i\in{\mathbb{Z}}$. A word $w=w(X)$ is encoded by a sequence of integers.
2.2. Subgroup graph
Here we review the definition of subgroup graphs and recall their basic properties. We assume the reader is familiar with this material and omit the proofs. All relevant proofs can be found in [7].
An $X$-digraph $\Gamma$ is a tuple $(V,E^{\pm},\mu,r)$, where
  • $(V,E^{\pm})$ defines a directed graph,
  • $r\in V$ is a designated vertex, called the root,
  • $\mu:E^{\pm}\to X^{\pm}$ is an edge labeling function (we often use notation $u\stackrel{{\scriptstyle x}}{{\to}}v$ for an edge $e$ labeled with $\mu(e)=x\in X^{\pm}$ that starts at $u$ and leads to $v$).
Define
[…]
called the set of positive and negative edges respectively. Clearly, $E^{\pm}=E^{+}\sqcup E^{-}$. We say that edges $e_{1},e_{2}\in E^{\pm}$ are inverses of each other if
$e_{1}=u\stackrel{{\scriptstyle x}}{{\to}}v\ \mbox{ and }\ e_{2}=v\stackrel{{\scriptstyle x^{-1}}}{{\to}}u,$
i.e., if they have the same endpoints, opposite direction, and opposite labels, in which case we write $e_{2}=e_{1}^{-1}$ and $e_{1}=e_{2}^{-1}$. We say that the edges in $\Gamma$ are inversible if $\Gamma$ with every edge $e=u\stackrel{{\scriptstyle x}}{{\to}}v$ contains its inverse. We say that $\Gamma$ is folded if for every $v\in V$ and $x\in X^{\pm}$ there exists at most one edge starting from $v$ labeled with $x$.
For an edge $e=u\stackrel{{\scriptstyle x}}{{\to}}v$ we denote its origin $u$ by $o(e)$ and its terminus $v$ by $t(e)$. A path $p$ in $\Gamma$ is a sequence of edges $e_{1},\dots,e_{t}$ satisfying the following connectedness condition:
$t(e_{s})=o(e_{s+1}),$
for every $s=1,\dots,t-1$. The label $\mu(p)$ of a path $p$ is the word
$\mu(p)=\mu(e_{1})\dots\mu(e_{t})\ \in\ (X^{\pm})^{\ast}.$
We say that $p$ is reduced if it does not contain consecutive opposite edges $ee^{-1}$. To reduce $p$ means to delete all pairs of consecutive opposite edges from $p$. It is easy to show that the result of path-reduction is uniquely defined, i.e., it does not depend on the sequence of reductions.
A circuit in $\Gamma$ is a closed path from $r$ to $r$. We say that $\Gamma$ is a core graph if for every edge $e$ there exists a reduced circuit in $\Gamma$ containing $e$. An $X$-digraph $\Gamma=(V,E^{\pm},\mu,r)$ is called a subgroup graph if it is a core graph, is folded and connected, and has inversible edges.
If $\Gamma$ is not folded, then there are distinct edges $e_{1}=v\stackrel{{\scriptstyle x}}{{\to}}u_{1}$ and $e_{2}=v\stackrel{{\scriptstyle x}}{{\to}}u_{2}$ with the same origin $v$ and the same label $x$. Identifying the edges $e_{1}$ and $e_{2}$ (and vertices $u_{1}$ and $u_{2}$) defines a single folding step. A sequence of foldings eventually terminates with a folded graph because each folding step decreases the size of $\Gamma$. It can be shown that the result does not depend on the specific sequence of foldings applied. The folding can be performed in nearly linear time, see [14].
Folded graphs have the following important property: for any path $p$ we have
[…]
We say that an $X$-digraph $\Gamma=(V,E^{\pm},\mu,r)$ accepts a word $w\in F(X)$ if $\Gamma$ contains a path $p$ from $r$ to $r$ labeled with $w$. The language of all accepted words is defined by
(2) $L(\Gamma)=L(\Gamma,r)=\left\{\,w\in F(X)\;\middle|\;\Gamma\mbox{ accepts }w\,\right\}.$
It is easy to see that $L(\Gamma)$ is a subgroup of $F(X)$ when $\Gamma$ is a subgroup graph.
2.3. A basis for $L(\Gamma)$
Let $\Gamma=(V,E^{\pm},\mu,r)$ be a subgroup graph. In this section we outline a procedure for finding a free basis for the subgroup $L(\Gamma)$.
Since $\Gamma$ is inversible, the set $E^{+}$ uniquely defines the set $E^{-}$. Hence, we can regard each pair of edges $\{e,e^{-1}\}$ as a single edge traversable in both directions, reading the label $x$ going in one direction and $x^{-1}$ in the other. From this perspective $(V,E^{\pm})$ can be viewed as an undirected graph $(V,E)$, where the edges $E$ are uniquely defined by $E^{+}$. A path in $(V,E)$ is a sequence of edges $e_{1},\dots,e_{k}$ from $E^{+}$, where each edge is either traversed in the forward (direct) direction or in the inverse direction.
We say that $T\subseteq E^{+}$ defines a spanning tree in $\Gamma$ if $(V,T)$ is a tree as an undirected graph. For a vertex $v\in V$ let $r,v]_{T}$ be the unique reduced path in $T$ from $r$ to $v$ and $\mu([r,v]_{T})$ its label. For $e=u\stackrel{{\scriptstyle x}}{{\to}}v\in E^{+}$ define the circuit
(3) $p_{e}=[r,o(e)]_{T}\cdot e\cdot[t(e),r]_{T}$
from $r$ to $r$ in $\Gamma$ and its label $w_{e}=\mu(p_{e})$. Clearly, $w_{e}=1$ if and only if $e\in T$.
Proposition 2.1 ([7, Lemma 6.1]).
$L(\Gamma)={\left\langle w_{e}\mid e\in E^{+}\setminus T\right\rangle}$.
2.4. Schreier graph
Recall that a right coset of a subgroup $H\leq G$ is the set
$Hg=\left\{\,hg\;\middle|\;h\in H\,\right\}.$
The collection of right cosets forms a partition of $G$. The number of distinct cosets of $H$ in $G$ is called the index of $H$ in $G$, denoted by $|G:H|$.
Consider a subgroup $H\leq G$ of a group $G$ generated by $x_{1},\dots,x_{n}\in G$. The Schreier graph of $H$ with respect to a generating set $X=\{x_{1},\dots,x_{n}\}$ is an $X$-digraph $\operatorname{{Sch}}(H,X)=(V,E,\mu,1_{H})$ defined by
$V=\left\{\,Hg\;\middle|\;g\in G\,\right\}\mbox{ and }E=\left\{\,Hg\stackrel{{\scriptstyle x}}{{\to}}Hgx\;\middle|\;g\in G,\ x\in X^{\pm}\,\right\},$
with the designated root $1_{H}=H\cdot 1\in V$, where $1$ is the identity in $G$. By construction, $\operatorname{{Sch}}(H,X)$ is
  • folded and connected;
  • has inversible edges;
  • in general, it is not a core graph;
  • $|V|=[G:H]$;
  • $L(\Gamma)=H$.
For an $X$-digraph $\Gamma$ and $v\in V(\Gamma)$ define the core $\operatorname{{Core}}(\Gamma,v)$ of $\Gamma$ with respect to $v$ as the subgraph induced by all reduced paths from $v$ to $v$ in $\Gamma$. It is easy to see that $\Gamma^{\prime}=\operatorname{{Core}}(\Gamma,r)$, where $\Gamma=(V,E,\mu,r)$, is a core graph defining the same subgroup, i.e., $L(\Gamma)=L(\Gamma^{\prime})$.
Theorem 2.2 ([7, Theorem 5.1, Theorem 5.2, and Definition 5.3]).
If $H$ is a subgroup of $F(X)$, then there is a unique (up to an isomorphism) subgroup graph $\Gamma$ satisfying $L(\Gamma)=H$. Denote such graph by $\Gamma_{H}$.
Proof.
In fact, $\operatorname{{Core}}(\operatorname{{Sch}}(H,X))$ is the required graph. ∎
2.5. Subgroup graph homomorphism
Let $\Gamma_{i}=(V_{i},E_{i},\mu_{i},r_{i})$ for $i=1,2$ be subgroup graphs. Recall that a map $\varphi:V_{1}\to V_{2}$ is a subgroup graph homomorphism if
  • $\varphi(r_{1})=r_{2}$;
[…]
Proposition 2.3 ([7, Lemma 4.1 and Proposition 4.3]).
$H_{1}\leq H_{2}\ \Leftrightarrow\$there exists a homomorphism $\varphi:\Gamma_{H_{2}}\to\Gamma_{H_{1}}$.
2.6. Regularity, self-similarity, and shift operation
Here we introduce the shift operation on subgroup graphs and discuss two properties that allow it to be computed efficiently. We say that a subgroup graph $\Gamma=(V,E^{\pm},\mu,r)$ is $X$-regular (or deterministic) if for each vertex $v$ of $\Gamma$ and for each $x\in X^{\pm}$, there is exactly one edge from $v$ labeled with $x$.
Proposition 2.4 ([7, cf. Proposition 8.3]).
Let $\Gamma$ be the subgroup graph of $H\leq F(X)$. Then
$F:H]=\begin{cases}|\Gamma|&\mbox{ if }\Gamma\mbox{ is $X$-regular,}\\ \infty&\mbox{ otherwise.}\end{cases}$
In particular, $F:H]<\infty\ \ \Leftrightarrow\ \ \Gamma_{H}$ is finite and $X$-regular.
By $\operatorname{{Aut}}(\Gamma)$ we denote the group of automorphisms of $\Gamma$. We say that $\Gamma=(V,E^{\pm},\mu,r)$ is self-similar if for every $u,v\in V$ there is an automorphism of $\Gamma$ that maps $u$ to $v$; such automorphism is unique when exists. Denote that automorphism by $S_{u,v}$. Note that $S_{u,v}$ induces
  • a permutation on the set of vertices $V$;
  • a permutation on the set of edges $E^{\pm}$;
  • a bijection from sequences of edges to sequences of edges
    $e_{1}\dots e_{k}\ \ \stackrel{{\scriptstyle S_{u,v}}}{{\mapsto}}\ \ S_{u,v}(e_{1})\dots S_{u,v}(e_{k}),$
    and the corresponding bijection from the set of paths that start at the vertex $u$ to the set of paths that start at $v$.
We use the same notation $S_{u,v}$ for the induced functions. Clearly, shift operators preserve labels, i.e., for every $u,v$ and a sequence of edges $p$ we have
$\mu(S_{u,v}(p))=\mu(p).$
Theorem 2.5 ([7, Theorem 8.14]).
$H\unlhd F(X)$ if and only if $\Gamma_{H}$ is $X$-regular and self-similar.
3. Preliminaries: straight-line programs
In this section we review the definition of straight-line programs, following the exposition in [11, Chapter 19]. See also [8] for further background.
3.1. Definition of a straight-line program
Formally, a straight-line program (SLP) is a quadruple $P=(X,{\mathcal{N}},R,\delta)$, where
  • $X=\{x_{1},\ldots,x_{n},\varepsilon\}$ is a finite set of terminal symbols (the alphabet), where $\varepsilon$ is a special symbol that denotes the empty string.
  • $\mathcal{N}$} is a finite set of non-terminal symbols.
  • $R\in{\mathcal{N}}$ is the root symbol.
  • $\delta:{\mathcal{N}}\to X\cup({\mathcal{N}}\times{\mathcal{N}})$ is a production function that determines the set of production rules. There are two types of production rules defined by $\delta(N)$ for $N\in{\mathcal{N}}$:
    • $\delta(N)=x\in X$,
    • $\delta(N)=(A,B)\in{\mathcal{N}}\times{\mathcal{N}}$.
    To be called an SLP, $P$ must define an acyclic production graph, defined below.
The production graph for $P$ is a directed graph $G(P)=(V,E)$, where $V=X\sqcup{\mathcal{N}}$ and
$E=$ $\ \ \ \{(N,\delta(N))\mid\delta(N)\in X\}$
$\cup\{(N,A)\mid\delta(N)=(A,B)\mbox{ for some }B\in{\mathcal{N}}\}$
$\cup\{(N,B)\mid\delta(N)=(A,B)\mbox{ for some }A\in{\mathcal{N}}\}.$
The graph $G(P)$ is acyclic if it does not contain a directed cycle.
For an SLP $P=(X,{\mathcal{N}},R,\delta)$ inductively define a function $\operatorname{val}:{\mathcal{N}}\rightarrow X^{\ast}$ by
[…]
and the sequence $\operatorname{val}(P)$ as $\operatorname{val}(R).$ The word $\operatorname{val}(P)$ is called the output of $P$. If $X$ is a group alphabet and $\operatorname{val}(N)$ is a reduced word for every $N\in{\mathcal{N}}$, then we say that $P$ is reduced.
In all cases considered in this paper the set of terminals $X$ is fixed. Therefore, we define the size of an SLP $P$ as the size of $\mathcal{N}$}, denoted by $|P|$.
3.2. Data representation for SLPs
In all our computations the alphabet $X$ is fixed and all operations on SLPs are actually performed on $\mathcal{N}$} and $\delta$. To simplify analysis, we make two assumptions.
  • (Assumption-I). We have a sufficiently large pool of symbols available for non-terminals and that it takes $O(1)$ time to generate a symbol not involved in any of the currently used SLPs.
  • (Assumption-II). The function $\delta$ is stored in a container that enables constant $O(1)$ time complexity for the following manipulations:
    • for a given $N\in{\mathcal{N}}$ get $\delta(N)$;
    • for a given $N\in{\mathcal{N}}$ delete the production for $N$;
    • for a given $N\in{\mathcal{N}}$ and $pr\in X\cup({\mathcal{N}}\times{\mathcal{N}})$ add the production $\delta(N)=pr$ to $\delta$;
    • for a given $N\in{\mathcal{N}}$ modify the value of $\delta(N)$.
    In particular, for two functions $\delta_{1},\delta_{2}$ with disjoint supports there is a procedure that adds the description of $\delta_{2}$ to the description of $\delta_{1}$ in $O(|\delta_{2}|)$ time.
We claim that we can make these assumptions when analyzing polynomial-time complexity. Indeed, let us compare our assumptions to a more realistic implementation for SLPs that defines a non-terminal as a natural number and $\delta$ as a trie that maps natural numbers from $\mathcal{N}$$\subset{\mathbb{Z}}$ (written in binary) to elements in $X\cup({\mathcal{N}}\times{\mathcal{N}})$. For this implementation, the $O(1)$ constant-time bound is replaced by an $O(\log_{2}(|{\mathcal{N}}|))$ bound, and the overall time complexity increases by a factor of $\log_{2}(|{\mathcal{N}}^{\ast}|)$, where $|{\mathcal{N}}^{\ast}|$ denotes the size of the largest SLP used in the computations. Clearly, this choice of defining $\delta$ as a trie preserves the property to be polynomial-time computable.
3.3. Basic properties
Here we discuss basic computational properties of SLPs.
Lemma 3.1.
For a given SLP $P$ it takes $O(|P|)$ time to decide whether $\operatorname{val}(P)=\varepsilon$.
Proof.
Clearly, for any $N\in{\mathcal{N}}$ we have $\operatorname{val}(N)=\varepsilon$ if and only if one of the following two conditions is satisfied:
  • $\delta(N)=\varepsilon$, or
  • $\delta(N)=(A,B)$ and $\operatorname{val}(A)=\varepsilon$ and $\operatorname{val}(B)=\varepsilon$.
Hence, we can decide if $\operatorname{val}(N)=\varepsilon$ for all non-terminals $N\in{\mathcal{N}}$ in linear time $O(|{\mathcal{N}}|)$ starting from the root $R$. ∎
For $N\in{\mathcal{N}}$ denote by $\operatorname{{first}}(N)$ and $\operatorname{{last}}(N)$ the first and the last element in $\operatorname{val}(N)$ respectively, if $\operatorname{val}(N)\neq\varepsilon$. If $\operatorname{val}(N)=\varepsilon$, then we write $\operatorname{{first}}(N)=\operatorname{{last}}(N)=\varnothing$.
Lemma 3.2.
Given an SLP $P=(X,{\mathcal{N}},R,\delta)$, it takes $O(|P|)$ time to compute $\operatorname{{first}}(R)$ and $\operatorname{{last}}(R)$.
Proof.
Clearly, for every $N\in{\mathcal{N}}$ we have
  • $\operatorname{{first}}(N)=\varnothing$ if $\delta(N)=\varepsilon$.
  • $\operatorname{{first}}(N)=x$ if $\delta(N)=x\in X$.
  • $\operatorname{{first}}(N)=\operatorname{{first}}(A)$ if $\delta(N)=(A,B)$ and $\operatorname{val}(A)\neq\varepsilon$.
  • $\operatorname{{first}}(N)=\operatorname{{first}}(B)$ if $\delta(N)=(A,B)$ and $\operatorname{val}(A)=\varepsilon$.
Clearly, we can use these formulae to compute $\operatorname{{first}}(N)$ for all non-terminals $n\in{\mathcal{N}}$ in linear time $O(|{\mathcal{N}}|)$ starting from the root $R$. $\operatorname{{last}}(N)$ can be computed in a similar way. ∎
Lemma 3.3.
For a given word $w=x_{1}\dots x_{k}$, where $x_{i}\in X$, it requires $O(|w|)$ time to construct an SLP $P_{w}$ satisfying $\operatorname{val}(P_{w})=w$.
Proof.
Clearly, the statement holds when $|w|=0$ or $|w|=1$. Let $X^{\prime}\subseteq X$ be the set of letters involved in $w$. Let ${\mathcal{N}}=X^{\prime}\cup\{A_{1},\dots,A_{k-1}\}$. Define an SLP $P_{w}=(X,{\mathcal{N}},A_{k},\delta)$, where $\delta$ is defined as follows:
  • $\delta(A_{x})=x$ for $x\in X^{\prime}$,
  • $\delta(A_{1})=(A_{x_{1}},A_{x_{2}})$,
  • $\delta(A_{2})=(A_{1},A_{x_{3}}),\dots,\delta(A_{k-1})=(A_{k-1},A_{x_{k}})$.
$P_{w}$ can be constructed in $O(|w|)$ time and satisfies $\operatorname{val}(P_{w})=w$. ∎
3.4. SLP concatenation
Consider two straight-line programs $P_{1}=(X,{\mathcal{N}}_{1},R_{1},\delta_{1})$ and $P_{2}=(X,{\mathcal{N}}_{2},R_{2},\delta_{2})$ over the same alphabet $X$. Assuming that ${\mathcal{N}}_{1}\cap{\mathcal{N}}_{2}=\varnothing$ and $A\notin{\mathcal{N}}_{1}\cup{\mathcal{N}}_{2}$ define a new SLP $P=(X,{\mathcal{N}}_{1}\cup{\mathcal{N}}_{2}\cup\{A\},A,\delta)$, where $\delta$ is defined by
$\delta(N)=\begin{cases}\delta_{1}(N)&\mbox{ if }N\in{\mathcal{N}}_{1},\\ \delta_{2}(N)&\mbox{ if }N\in{\mathcal{N}}_{2},\\ (R_{1},R_{2})&\mbox{ if }N=A.\end{cases}$
Lemma 3.4.
$\operatorname{val}(P)=\operatorname{val}(P_{1})\circ\operatorname{val}(P_{2})$.
Proof.
$\operatorname{val}(P)=\operatorname{val}(A)=\operatorname{val}(R_{1})\circ\operatorname{val}(R_{2})=\operatorname{val}(P_{1})\circ\operatorname{val}(P_{2})$. ∎
Denote the SLP $P$ by $P_{1}\circ P_{2}$. More generally, for SLPs $P_{1},\dots,P_{k}$ denote by $P_{1}\circ\dots\circ P_{k}$ the SLP $((P_{1}\circ P_{2})\circ P_{3})\dots\circ P_{k}$. Notice that concatenating $k$ SLPs requires k1k-1 additional non-terminals.
3.5. Straight-line program over an $X$-digraph
Let $\Gamma=(V,E^{\pm},\mu,r)$ be a subgroup graph over the alphabet $X$. We can treat the set of edges $E^{\pm}$ as an alphabet. Note that $E^{\pm}$ forms a group alphabet since, by assumption, $\Gamma$ contains with every edge $e$ its inverse $e^{-1}$. Hence, we can work with SLPs over $E^{\pm}$. The output $\operatorname{val}(P)$ of such SLP $P$ is a sequence of edges in $\Gamma$.
Let $P$ be an SLP over an $X$-digraph $\Gamma$. For $N\in{\mathcal{N}}$ define vertices $o(N)$ and $t(N)$ as
$o(N)=o(\operatorname{{first}}(N))\mbox{ and }t(N)=t(\operatorname{{last}}(N)),$
if $\operatorname{val}(N)\neq\varepsilon$ and as $\varnothing$ if $\operatorname{val}(N)=\varepsilon$.
Lemma 3.5.
It takes linear time to decide if the sequence of edges $\operatorname{val}(P)$ is a path.
Proof.
Clearly, $\operatorname{val}(P)$ is a path if and only if
$\forall N\in{\mathcal{N}}\ \ \ \delta(N)=(A,B)\ \wedge\ \operatorname{val}(A)\neq\varepsilon\ \wedge\ \operatorname{val}(B)\neq\varepsilon\ \rightarrow\ t(A)=o(B).$
This condition can be checked in linear time because, by Lemma 3.2, $t(A)$ and $o(B)$ can be computed in linear time for all non-terminals in $P$. ∎
In the next proposition we assume that $\Gamma$ is a fixed subgroup graph. This allows us to treat all relevant data related to $\Gamma$, such as an explicit description of the shift operation $S_{u,v}$, as precomputed.
Proposition 3.6.
Let $\Gamma=(V,E^{\pm},\mu,r)$ be an $X$-regular and self-similar subgroup graph. Let $u,v\in V$ and $S_{u,v}:E^{\pm}\to E^{\pm}$ be a permutation on the set of edges given explicitly as a set of pairs $(e,S_{u,v}(e))$. Given an SLP $P$ over $\Gamma$ it requires $O(|P|)$ time to construct an SLP $P^{\prime}$ satisfying
  • $\operatorname{val}(P^{\prime})=S_{u,v}(\operatorname{val}(P))$,
  • $|P^{\prime}|=|P|$.
Proof.
For every non-terminal $N$ such that $\delta(N)=e\in E^{\pm}$, the procedure replaces $e$ with $S_{u,v}(e)$. This does not change the number of non-terminals. ∎
Denote by $S_{u,v}(P)$ the SLP constructed in the proof of Proposition 3.6 for $P$.
We say that $P$ is reduced if its label $\operatorname{val}(P)$ is reduced as an element of $F(E)$. To reduce $P$ means to find an SLP $P^{\prime}$ such that $o(P^{\prime})=o(P)$, $\mu(\operatorname{val}(P^{\prime}))=\mu(\operatorname{val}(P))$, and $\operatorname{val}(P^{\prime})$ is reduced.
Theorem 3.7 ([9], Theorem 4.5).
It takes polynomial-time to reduce $P$.
4. The case of equal associated normal subgroups of finite index
In this section, we assume that the associated subgroups $H_{1}$ and $H_{2}$ are equal, normal in $F$, and of finite index. Let $\Gamma=(V,E^{\pm},\mu,r)$ be the subgroup graph for $H_{1}=H_{2}$. These assumptions imply that
  • $\Gamma$ is finite.
  • $\Gamma$ is $X$-regular.
  • For any $u,v\in V$ there is an automorphism $\varphi_{u,v}:\Gamma\to\Gamma$ satisfying $\varphi_{u,v}(u)=v$.
Since the group $G$ is fixed, we treat the following data as part of its description.
  • The subgroup graph $\Gamma=(V,E^{\pm},r,\mu)$ for $H_{1}=H_{2}$.
  • A set of edges $T\subseteq E^{+}$ defining a spanning tree in $\Gamma$ as described in Section 2.3.
  • For every $e\in E^{+}\setminus T$ we have
    • the circuit $p_{e}$ in $\Gamma$ corresponding to $e$ defined by (3);
    • the circuit $p_{e}^{\varphi}$ in $\Gamma$ satisfying $\mu(p_{e}^{\varphi})=\varphi(\mu(p_{e}))$;
    • the circuit $p_{e}^{\varphi^{-1}}$ in $\Gamma$ satisfying $\mu(p_{e}^{\varphi^{-1}})=\varphi^{-1}(\mu(p_{e}))$;
    • an SLP $P_{e}^{\varphi}$ satisfying $\operatorname{val}(P_{e}^{\varphi})=p_{e}^{\varphi}$;
    • an SLP $P_{e^{-1}}^{\varphi}$ satisfying $\operatorname{val}(P_{e^{-1}}^{\varphi})=(p_{e}^{\varphi})^{-1}$;
    • an SLP $P_{e}^{\varphi^{-1}}$ satisfying $\operatorname{val}(P_{e}^{\varphi^{-1}})=p_{e}^{\varphi^{-1}}$;
    • an SLP $P_{e^{-1}}^{\varphi^{-1}}$ satisfying $\operatorname{val}(P_{e^{-1}}^{\varphi^{-1}})=(p_{e}^{\varphi^{-1}})^{-1}$.
Define a constant
(4) $C=C_{\varphi}=\sum_{e\in E^{+}\setminus T}|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|+|P_{e}^{\varphi^{-1}}|+|P_{e^{-1}}^{\varphi^{-1}}|+2.$
Now we describe the algorithm for the word problem in $F\ast_{\varphi}t$. First, a given word (1) is translated into an alternating sequence
(5) $P_{0},t^{\varepsilon_{1}},P_{1},\dots,P_{k-1},t^{\varepsilon_{k}},P_{k}$
of letters $t^{\pm 1}$ and straight-line programs $P_{i}$ over the alphabet $E^{\pm}$ (a formal alphabet of edges of $\Gamma$) satisfying the following conditions:
  • $\operatorname{val}(P_{i})$ is a path in $\Gamma$ starting from $r$,
  • $\mu(\operatorname{val}(P_{i}))=w_{i}$,
using Lemma 3.3. By Lemma 3.3, (5) can be computed in linear time. All further computations are performed on the sequence (5).
4.1. Application of $\varphi^{\pm 1}$ to an SLP
Let $P$ be an SLP over $\Gamma$ such that $\operatorname{val}(P)$ is a circuit in $\Gamma$ from $r$ to $r$. Then $\mu(\operatorname{val}(P))\in H$ and $\varphi$ is applicable to $\mu(\operatorname{val}(P))$.
Proposition 4.1.
There is an algorithm that for a given $P$ produces an SLP $P^{\prime}$ in $O(|P|)$ time satisfying the following conditions:
  • (a)
    $\operatorname{val}(P^{\prime})$ is a circuit based at $r$;
  • (b)
    $\mu(\operatorname{val}(P^{\prime}))=\varphi(\mu(\operatorname{val}(P)))$;
  • (c)
    $|P^{\prime}|\leq|P|+\sum_{e\in E^{+}\setminus T}|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|$.
The same holds for $\varphi^{-1}$.
Proof.
The algorithm modifies the terminals $e\in E^{\pm}$ in $P$. It distinguishes two types of terminals.
(Case-I) For each terminal $e\in T$ the algorithm deletes $e$ and $e^{-1}$ from the definitions of all non-terminals. That effectively deletes all occurrences of $e$ and $e^{-1}$ from $\operatorname{val}(P)$.
(Case-II) For each terminal $e\in E^{+}\setminus T$ perform the following.
  • Add the description of $P_{e}^{\varphi}$ and $P_{e^{-1}}^{\varphi}$ to $P$.
  • Add non-terminals $N_{e}$ and $N_{e^{-1}}$ with productions $\delta(N_{e})=P_{e}^{\varphi}$ and $\delta(N_{e^{-1}})=P_{e^{-1}}^{\varphi}$.
  • Delete every non-terminal $N$ with production $\delta(N)=e$ from the description of $P$, and in all other productions replace each occurrence of $N$ with $N_{e}$.
  • Delete every non-terminal $N$ with production $\delta(N)=e^{-1}$ from the description of $P$, and in all other productions replace each occurrence of $N$ with $N_{e^{-1}}$.
This effectively replaces every occurrence of $e$ and $e^{-1}$ in $\operatorname{val}(P)$ with $\operatorname{val}(P_{e}^{\varphi})$ and $\operatorname{val}(P_{e^{-1}}^{\varphi})$, respectively.
By construction, all three properties hold for the obtained $P^{\prime}$. ∎
4.2. The word problem algorithm
Proposition 4.2.
Consider a segment $P_{i-1},t^{-1},P_{i},t,P_{i+1}$ in (5). It requires $O(|P_{i-1}|+|P_{i}|+|P_{i+1}|)$ time to check if $P_{i}$ defines an element in $H_{1}$ (i.e., if $\mu(\operatorname{val}(P_{i}))\in H_{1}$) and, if so, to construct an SLP $P$ satisfying the following properties:
  • (a)
    $\operatorname{val}(P)$ is a path in $\Gamma$ that starts at $r$,
  • (b)
    $\mu(\operatorname{val}(P))\ =_{G}\ \mu(\operatorname{val}(P_{i-1}))\cdot t^{-1}\mu(\operatorname{val}(P_{i}))t\cdot\mu(\operatorname{val}(P_{i+1}))$,
  • (c)
    $|P|\leq|P_{i-1}|+|P_{i}|+|P_{i+1}|+C$.
The same holds for segments $P_{i-1},t,P_{i},t^{-1},P_{i+1}$, when $P_{i}$ defines an element in $H_{2}$.
Proof.
By Lemmas 3.2 and 3.5, we can check if $\operatorname{val}(P_{i})$ defines a path that starts and ends at $r$ (i.e., if $\mu(\operatorname{val}(P_{i}))\in H_{1}$), in linear $O(|P_{i}|)$ time. By Proposition 4.1, in linear $O(|P_{i}|)$ time we can compute an SLP $P_{i}^{\prime}$ satisfying
  • $\operatorname{val}(P_{i}^{\prime})$ a circuit in $\Gamma$ from $r$ to $r$, and
  • $\mu(\operatorname{val}(P_{i}^{\prime}))=_{G}t^{-1}\mu(\operatorname{val}(P_{i}))t$.
Hence, the word
$\mu(\operatorname{val}(P_{i-1}))\mu(\operatorname{val}(P_{i}^{\prime}))\mu(\operatorname{val}(P_{i+1}))$
defines the same element as the right-hand side of (b). Note that, in general, $\operatorname{val}(P_{i-1})\circ\operatorname{val}(P_{i}^{\prime})\circ\operatorname{val}(P_{i+1}))$ does not define a continuous path in $\Gamma$; there may be up to two points of discontinuity.
To create a required SLP $P$, it remains to properly concatenate the paths $\operatorname{val}(P_{i-1})$, $\operatorname{val}(P_{i}^{\prime})$, $\operatorname{val}(P_{i+1})$, which can be done using shift operators. Use Lemma 3.2 and Proposition 3.6 to compute
$v_{1}=t(\operatorname{{last}}(P_{i-1})),\ P_{i}^{\prime\prime}=S_{r,v_{1}}(P_{i}^{\prime})\ \mbox{ and }\ v_{2}=t(\operatorname{{last}}(P_{i}^{\prime\prime})),\ P_{i+1}^{\prime}=S_{r,v_{2}}(P_{i+1}).$
By definition of $P_{i}^{\prime\prime}$ and $P_{i+1}^{\prime}$, concatenation $\operatorname{val}(P_{i-1})\circ\operatorname{val}(P_{i}^{\prime\prime})\circ\operatorname{val}(P_{i+1}^{\prime})$ is a path in $\Gamma$ that starts at $r$. Its label defines the same element as the right-hand side of (b) because shift operators preserve labels. Therefore, the SLP $P=P_{i-1}\circ P_{i}^{\prime\prime}\circ P_{i+1}^{\prime}$ satisfies (a) and (b).
Finally, by construction,
$|P_{i}^{\prime\prime}|\leq|P_{i}^{\prime}|\leq|P_{i}|+\sum_{e\in E^{+}\setminus T}|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|\ \mbox{ and }\ |P_{i+1}^{\prime}|\leq|P_{i+1}|$
because applying shift operators does not increase the size of an SLP. Thus,
$|P|\leq|P_{i-1}|+|P_{i}|+|P_{i+1}|+\sum_{e\in E^{+}\setminus T}|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|+2\leq|P_{i-1}|+|P_{i}|+|P_{i+1}|+C$
and $P$ satisfies (c). ∎
Theorem 4.3.
Suppose that $H_{1}=H_{2}$ are normal subgroups of $F$ of finite index and let $\varphi:H_{1}\to H_{2}$ be an isomorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
Proof.
Consider a word $w$ of type (1). If the syllable length $k$ of $w$ is trivial, then we directly check if $w_{0}$ is trivial in the base group $F$.
Suppose that $k\geq 1$. Translate $w$ into a sequence (5), which can be done in $O(|w|)$ time. Then apply a sequence of Britton’s reductions using Proposition 4.2 for a single reduction step. If $w=_{G}1$, then the process produces a single SLP $P^{\ast}$ satisfying
$|P^{\ast}|\leq\sum_{i=0}^{k}|P_{i}|+\tfrac{k}{2}C\leq(C+1)\sum_{i=0}^{k}|P_{i}|$
which is $O(|w|)$ because $C$ is a fixed parameter of the group. The time complexity of reducing (5) to $P^{\ast}$ can be bounded by $O(k|w|)$ or simply $O(|w|^{2})$. Finally, it remains to check if $\mu(\operatorname{val}(P^{\ast}))=\varepsilon$ in $F$. By Theorem 3.7, that can be done in polynomial time. Therefore, the total time-complexity of the described procedure can be bounded by a polynomial. ∎
5. The case of equal associated subgroups of finite index
In this section, we generalize the algorithm from Section 4.2 to the case when $\varphi:H\to H$ can be restricted to a normal subgroup of $F$ of finite index, in which case we say that $\varphi$ is normalizable.
5.1. $\varphi$-stable subgroups
Let $H\leq F$ and $\varphi\in\operatorname{{Aut}}(H)$. We say that $H^{\prime}\leq H$ is $\varphi$-stable if $\varphi(H^{\prime})\subseteq H^{\prime}$, i.e., if $\varphi|_{H^{\prime}}\in\operatorname{{Aut}}(H^{\prime})$.
Lemma 5.1 (Join).
If $H^{\prime},H^{\prime\prime}$ are $\varphi$-stable, then $\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle$} is $\varphi$-stable.
Proof.
For any $g\in H^{\prime}\cup H^{\prime\prime}$ we have $\varphi(g)\in\varphi(H^{\prime})\cup\varphi(H^{\prime\prime})=H^{\prime}\cup H^{\prime\prime}$. Hence, $\varphi(H^{\prime}\cup H^{\prime\prime})\subseteq H^{\prime}\cup H^{\prime\prime}$. Therefore, $\varphi({\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle})\subseteq{\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle}$. ∎
Lemma 5.2 (Meet).
If $H^{\prime},H^{\prime\prime}$ are $\varphi$-stable, then $H^{\prime}\cap H^{\prime\prime}$ is $\varphi$-stable.
Proof.
For any $g\in H^{\prime}\cap H^{\prime\prime}$ we have $\varphi(g)\in\varphi(H^{\prime})\cap\varphi(H^{\prime\prime})=H^{\prime}\cap H^{\prime\prime}$. Hence, $\varphi(H^{\prime}\cap H^{\prime\prime})\subseteq H^{\prime}\cap H^{\prime\prime}$. ∎
Thus, the set of all $\varphi$-stable subgroups, denoted by $L_{\varphi}$, has a structure of a bounded lattice, with the maximum element $H$ and the minimum element $\{1\}$. Let us consider a set
$L_{\varphi}^{\ast}=\left\{\,N\in L_{\varphi}\;\middle|\;N\unlhd F\,\right\}.$
$L_{\varphi}^{\ast}$ is not empty, as it contains $\{1\}$. Furthermore, it is easy to check that it is a sublattice of $L_{\varphi}$. Denote by $M_{\varphi}$ the maximum element of $L_{\varphi}^{\ast}$.
5.2. $\varphi$-stable normal subgroups of finite index
Now, let $H\leq F$ be a subgroup of finite index and $\varphi\in\operatorname{{Aut}}(H)$. We say that $\varphi$ is normalizable if $\varphi$ can be restricted to a normal subgroup $N\unlhd F$ of finite index (i.e., if $M_{\varphi}$ has finite index).
Problem 5.3.
Is it true that every $\varphi$ is normalizable?
We suspect that the answer is negative in general. However, how can one find such a subgroup $N$ if it exists? Let us review the properties of a required subgroup $N$. It should satisfy the following four properties:
  • NN is normal,
  • NN has finite index,
  • $c^{-1}Nc\subseteq N$ for every $c\in F$,
  • $\varphi^{\pm 1}(N)\subseteq N$.
Let us define the following sequence of subgroups:
(6) $\begin{array}[]{l}H_{0}=H,\\ H_{i+1}=\cap_{c}c^{-1}H_{i}c\cap\varphi(H_{i})\cap\varphi^{-1}(H_{i})\ \ \ \mbox{ for }i\geq 0.\end{array}$
Lemma 5.4.
$F:H_{k}]<\infty$ for every $k\geq 0$.
Proof.
Induction on $k$. By assumption, $H_{0}=H$ has finite index. Since, $\varphi^{\pm 1}(H)=H$ we have
$H_{1}=\cap_{c}c^{-1}H_{0}c$
which has finite index in $F$. Assume that the statement holds for $H_{k-1}$ and consider $H_{k}$.
$[F:H_{k-1}]<\infty$ $\ \ \Rightarrow\ \ [H:H_{k-1}]<\infty$
$\ \ \Rightarrow\ \ [H:\varphi(H_{k-1})]<\infty$ (because $\varphi\in\operatorname{{Aut}}(H)$)
$\ \ \Rightarrow\ \ [F:\varphi(H_{k-1})]<\infty.$
Similarly $F:\varphi^{-1}(H_{k-1})]<\infty$. Hence, $H_{k}$ is an intersection of finitely many subgroups of finite index and, hence, has finite index itself. ∎
Lemma 5.5.
$M_{\varphi}\leq H_{k}$ for every $k\geq 0$.
Proof.
Induction on $k$. The statement holds for $k=0$. Assume that the statement holds for k1k-1, i.e., $M_{\varphi}\leq H_{k-1}$. Then the following holds.
$M_{\varphi}$ $=\cap_{c}c^{-1}M_{\varphi}c\leq\cap_{c}c^{-1}H_{k-1}c.$
$M_{\varphi}$ $=\varphi(M_{\varphi})\leq\varphi(H_{k-1}).$
$M_{\varphi}$ $=\varphi^{-1}(M_{\varphi})\leq\varphi^{-1}(H_{k-1}).$
Therefore, $M_{\varphi}\leq\cap_{c}c^{-1}H_{k-1}c\cap\varphi(H_{k-1})\cap\varphi^{-1}(H_{k-1})=H_{k}$. ∎
Lemma 5.6.
$\cap_{i}H_{i}=M_{\varphi}$.
Proof.
Denote $\cap_{i}H_{i}$ by $L$. It is easy to see that it is a normal subgroup of $F$. Also
$x\in L$ $\ \ \Rightarrow\ \ x\in H_{i}\ \forall i$
$\ \ \Rightarrow\ \ \varphi(x)\in\varphi(H_{i})\subseteq H_{i+1}\ \forall i$
$\ \ \Rightarrow\ \ \varphi(x)\in L.$
Hence, $L\in L_{\varphi}^{\ast}$. By Lemma 5.5, $M_{\varphi}\leq L$. Thus, $L=M_{\varphi}$. ∎
We say that the sequence (6) stabilizes if $H_{i+1}=H_{i}$ for some index $i$. The next lemma follows from the definition of $H_{i+1}$.
Lemma 5.7.
If $H_{i+1}=H_{i}$, then $H_{i}=M_{\varphi}$.
Proposition 5.8.
$\varphi$ is normalizable if and only if $\{H_{i}\}$ stabilizes.
Proof.
$\Leftarrow$” Lemma 5.7.
$\Rightarrow$” If $F:M_{\varphi}]=m<\infty$, then $H_{\lceil\log_{2}(m)\rceil}=M_{\varphi}$ because $H_{i}:H_{i+1}]\geq 2$ whenever $H_{i+1}\neq H_{i}$. ∎
5.3. The case when $\varphi$ extends to an automorphism of $F$
Recall that a subgroup $N\leq F$ is characteristic if $\varphi(N)\subseteq N$ for every $\varphi\in\operatorname{{Aut}}(F)$. It is easy to see that every characteristic subgroup is normal.
Lemma 5.9.
Let $H$ be a finite index subgroup of $F$. Then $H$ contains a characteristic subgroup of $F$ of finite index.
Proof.
Let $F:H]=n<\infty$. Then $F:\varphi(H)]=n$ for any $\varphi\in\operatorname{{Aut}}(F)$. We know that for a finitely generated group, it has a finite number of subgroups of index $n$ for any $n\in{\mathbb{N}}$. The number of subgroups of index $n$ in $F$ is finite. Therefore
$N=\bigcap_{\varphi\in\operatorname{{Aut}}(F)}\varphi(H)$
has finite index, is characteristic, and is contained in $H$. ∎
Proposition 5.10.
If $\varphi\in\operatorname{{Aut}}(H)$ extends to an automorphism of $F$, then $\varphi$ is normalizable.
Proof.
By Lemma 5.9, $H$ contains a finite index characteristic subgroup $N$ of $F$. ∎
5.4. ${\mathbf{WP}}(F\ast_{\varphi}t)$ is decidable in polynomial time when $\varphi$ is normalizable
Suppose that $\varphi$ is normalizable and $N$ is a normal subgroup of $F$ of finite index satisfying $\varphi|_{N}\in\operatorname{{Aut}}(N)$. Denote $\varphi|_{N}$ by $\varphi^{\ast}$.
5.4.1. Syllables
Let $\Gamma^{\ast}=(V,E^{\pm},r,\mu)$ be the subgroup graph for $N$. Fix a set of edges $T\subseteq E^{+}$ defining a spanning tree in $\Gamma$ as described in Section 2.3. The graph $\Gamma^{\ast}$ is the Cayley graph of the finite group $F/N$, which induces a natural group operation on its vertex set $V$. Moreover, since $N\leq H\leq F$, it follows that $H/N\leq F/N$. Hence, for $v\in V$ we write $v\in H$ whenever the coset corresponding to $v$ belongs to $H/N$. In particular, we have $r\in H$.
As in Section 2.3, fix a spanning tree $T\subseteq E^{+}$ in $\Gamma^{\ast}$. As in Section 4, we work with paths over $\Gamma^{\ast}$. In this case, however, each path is represented by a pair $(P,v)$, where $P$ is an SLP over the alphabet $E^{\pm}$ and $v\in V$, subject to the following condition:
  • $\operatorname{val}(P)$ is a circuit in $\Gamma^{\ast}$ based at $r$.
We call such a pair a syllable. Each syllable defines the path
$p(P,v)=\operatorname{val}(P)[r,v]_{T}$
in $\Gamma^{\ast}$ that starts at $r$, ends at $v$, and which label is
$w(P,v)=\mu(\operatorname{val}(P))\mu([r,v]_{T}).$
5.4.2. Precomputed data
Since the group $G=F\ast_{\varphi}t$ is fixed, we assume that the following data is included in its description.
  • The subgroup graph $\Gamma$ for $H$.
  • The subgroup graph $\Gamma^{\ast}=(V,E^{\pm},r,\mu)$ for $N$.
  • A set of edges $T\subseteq E^{+}$ defining a spanning tree in $\Gamma$ as described in Section 2.3.
  • The index $k=[F:N]$ (the same as the order of $\Gamma^{\ast}$).
  • For every $e\in E^{+}\setminus T$ we have
    • the circuit $p_{e}$ in $\Gamma^{\ast}$ corresponding to $e$ defined by (3);
    • the circuit $p_{e}^{\varphi}$ in $\Gamma^{\ast}$ satisfying $\mu(p_{e}^{\varphi})=\varphi(\mu(p_{e}))$;
    • the circuit $p_{e}^{\varphi^{-1}}$ in $\Gamma^{\ast}$ satisfying $\mu(p_{e}^{\varphi^{-1}})=\varphi^{-1}(\mu(p_{e}))$;
    • an SLP $P_{e}^{\varphi}$ satisfying $\operatorname{val}(P_{e}^{\varphi})=p_{e}^{\varphi}$;
    • an SLP $P_{e^{-1}}^{\varphi}$ satisfying $\operatorname{val}(P_{e^{-1}}^{\varphi})=(p_{e}^{\varphi})^{-1}$;
    • an SLP $P_{e}^{\varphi^{-1}}$ satisfying $\operatorname{val}(P_{e}^{\varphi^{-1}})=p_{e}^{\varphi^{-1}}$;
    • an SLP $P_{e^{-1}}^{\varphi^{-1}}$ satisfying $\operatorname{val}(P_{e^{-1}}^{\varphi^{-1}})=(p_{e}^{\varphi^{-1}})^{-1}$.
  • For every vertex $v$ such that $v\in H$ we have a syllable $(P_{v}^{\varphi},v^{\prime})$ satisfying
    $w(P_{v}^{\varphi},v^{\prime})=\varphi(\mu([r,v]_{T})),$
    and a similar pair $(P_{v}^{\varphi^{-1}},v^{\prime})$ for $\varphi^{-1}$.
  • For every $v_{1},v_{2}\in V$ we have an SLP $T_{v_{1},v_{2}}$ constructed by Lemma 3.3 satisfying $\operatorname{val}(T_{v_{1},v_{2}})=[v_{1},v_{2}]_{T}$.
  • For every $v_{1},v_{2}\in V$ we have an SLP $C_{v_{1},v_{2}}=S_{r,v_{1}}(T_{r,v_{2}})\circ T_{v_{1}v_{2},r}$.
Define a constant
$C^{\ast}=$ $\sum_{e\in E^{+}\setminus T}(|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|+|P_{e}^{\varphi^{-1}}|+|P_{e^{-1}}^{\varphi^{-1}}|)$
$+\sum_{v\in V}(|P_{v}^{\varphi}|+|P_{v}^{\varphi^{-1}}|)$
$+2\sum_{v\in V}|T_{r,v}|+2\sum_{v_{1},v_{2}\in V}|C_{v_{1},v_{2}}|+7.$
5.4.3. Data representation for a given word $w$
A given word (1) is translated into an alternating sequence
(7) $(P_{0},v_{0}),t^{\varepsilon_{1}},(P_{1},v_{1}),\dots,(P_{k-1},v_{k-1}),t^{\varepsilon_{k}},(P_{k},v_{k})$
of letters $t^{\pm 1}$ and syllables $(P_{i},v_{i})$ satisfying $w(P_{i},v_{i})=w_{i}.$ The sequence (7) is computed by applying Lemma 5.11 to $w_{0},\dots,w_{k}$.
Lemma 5.11.
For a given $w\in F$, it requires $O(|w|)$ time to construct a syllable $(P,v)$ satisfying $w(P,v)=w$.
Proof.
Let $v$ be the endpoint of the path labeled by $w$ in $\Gamma^{\ast}$ starting from $r$. The word $w\cdot\mu([v,r]_{T})$ labels a circuit $c$ in $\Gamma^{\ast}$, and its length is bounded by $|w|+|\Gamma^{\ast}|$. Using Lemma 3.3, construct an SLP $P$ such that $\operatorname{val}(P)=c$. The pair $(P,v)$ is a required syllable. ∎
5.4.4. Application of $\varphi^{\pm 1}$ to $(P,v)$
Consider a syllable $(P,v)$. Obviously,
$w(P,v)\in H\ \ \Leftrightarrow\ \ v\in H.$
To apply $\varphi^{\pm}$ to $(P,v)$ means to compute a syllable $(P^{\prime},v^{\prime})$ satisfying $w(P^{\prime},v^{\prime})=\varphi^{\pm}(w(P,v))$.
Proposition 5.12.
There is an algorithm that for a given syllable $(P,v)$, that satisfies $w(P,v)\in H$, produces a syllable $(P^{\prime},v^{\prime})$ in $O(|P|)$ time satisfying the following conditions:
  • (a)
    $w(P^{\prime},v^{\prime})=\varphi(w(P,v))$,
  • (b)
    $|P^{\prime}|\leq|P|+\sum_{e\in E^{+}\setminus T}(|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|)+|P_{v}^{\varphi}|+1$.
A similar statement holds for $\varphi^{-1}$.
Proof.
Since $\mu(\operatorname{val}(P))\in N$, we can process $P$ using Proposition 4.1 and denote the result by $P_{1}$. That increases the number of non-terminals by at most $\sum_{e\in E^{+}\setminus T}(|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|)$. Then use the precomputed pair $(P_{v}^{\varphi},v^{\prime})$ defining $\varphi(\mu([r,v]_{T}))$ and concatenate $P_{1}$ and $P_{v}^{\varphi}$ to get $P^{\prime}$. That adds $|P_{v}^{\varphi}|+1$ non-terminals. Clearly, $(P^{\prime},v^{\prime})$ is a required syllable. ∎
5.4.5. Syllable concatenation
To concatenate syllables $(P_{1},v_{1})$ and $(P_{2},v_{2})$ means to construct a syllable $(P,v)$ satisfying
(8) $w(P_{1},v_{1})w(P_{2},v_{2})=w(P,v).$
One way to construct a required pair $(P,v)$ is to shift $p(P_{2},v_{2})$ in $\Gamma^{\ast}$ so that its origin is $v_{1}$ and attach the result to $(P_{1},v_{1})$, see Figure 1, and then construct a proper syllable.
Figure 1. Concatenation of the paths $p(P_{1},v_{1})$ and $p(P_{2},v_{2})$.
In the next proposition we prove that the syllable
(9) $(P,v)=(P_{1}\circ T_{r,v_{1}}\circ S_{r,v_{1}}(P_{2})\circ\underbrace{S_{r,v_{1}}(T_{r,v_{2}})\circ T_{v_{1}v_{2},r}}_{C_{v_{1},v_{2}}},v_{1}v_{2})$
defines concatenation of $(P_{1},v_{1})$ and $(P_{2},v_{2})$. We denote $(P,v)$ defined in (9) by $(P_{1},v_{1})\circ(P_{2},v_{2})$.
Proposition 5.13.
$(P,v)=(P_{1},v_{1})\circ(P_{2},v_{2})$ satisfies (8). Constructing $(P,v)$ requires $O(|P_{1}|+|P_{2}|)$ time. Moreover,
$|P|\leq|P_{1}|+|P_{2}|+|T_{r,v_{1}}|+|C_{v_{1},v_{2}}|+3.$
Proof.
By definition,
$w(P_{1},v_{1})=\mu(\operatorname{val}(P_{1}))\mu([r,v_{1}]_{T})$ $=\mu(\operatorname{val}(P_{1}))\mu(\operatorname{val}(T_{r,v_{1}})),$
$w(P_{2},v_{2})=\mu(\operatorname{val}(P_{2}))\mu([r,v_{2}]_{T})$ $=\mu(\operatorname{val}(P_{2}))\mu(\operatorname{val}(T_{r,v_{2}}))$
$=\mu(\operatorname{val}(S_{r,v_{1}}(P_{2})))\mu(S_{r,v_{1}}(T_{r,v_{2}})),$
because an application of $S_{r,v_{1}}$ does not change the labels. Hence,
$w(P_{1},v_{1})w(P_{2},v_{2})=\mu(\operatorname{val}(P_{1}\circ T_{r,v_{1}}\circ S_{r,v_{1}}(P_{2})\circ S_{r,v_{1}}(T_{r,v_{2}}))).$
Notice that $\operatorname{val}(P_{1}\circ T_{r,v_{1}}\circ S_{r,v_{1}}(P_{2})\circ S_{r,v_{1}}(T_{r,v_{2}}))$ is a path in $\Gamma^{\ast}$ from $r$ to $v_{1}v_{2}$. Since the path $\operatorname{val}(T_{v_{1}v_{2},r}\circ T_{r,v_{1}v_{2}})$ freely reduces to $\varepsilon$ as an element of $F(E)$, we have
$w(P_{1},v_{1})w(P_{2},v_{2})=\mu(\operatorname{val}(P_{1}\circ T_{r,v_{1}}\circ S_{r,v_{1}}(P_{2})\circ S_{r,v_{1}}(T_{r,v_{2}})\circ T_{v_{1}v_{2},r}\circ T_{r,v_{1}v_{2}})),$
where $\operatorname{val}(P_{1}\circ T_{r,v_{1}}\circ S_{r,v_{1}}(P_{2})\circ S_{r,v_{1}}(T_{r,v_{2}})\circ T_{v_{1}v_{2},r})$ is a circuit based at $r$ and $\operatorname{val}(T_{r,v_{1}v_{2}})=[r,v_{1}v_{2}]_{T}$. Thus, $(P_{1}\circ T_{r,v_{1}}\circ S_{r,v_{1}}(P_{2})\circ S_{r,v_{1}}(T_{r,v_{2}})\circ T_{v_{1}v_{2},r},v_{1}v_{2})$ satisfies (8).
We have $P=P_{1}\circ T_{r,v_{1}}\circ S_{r,v_{1}}(P_{2})\circ C_{v_{1},v_{2}}$, by definition of $C_{v_{1},v_{2}}$. By Proposition 3.6, $S_{r,v_{1}}(P_{2})$ can be computed in $O(|P_{2}|)$ time and satisfies $|S_{r,v_{1}}(P_{2})|=|P_{2}|$. Then concatenation $P_{1}\circ T_{r,v_{1}}\circ S_{r,v_{1}}(P_{2})\circ C_{v_{1},v_{2}}$ can be computed in $O(|P_{1}|+|P_{2}|)$ time because $T_{r,v_{1}}$ and $C_{v_{1},v_{2}}$ are of constant size. The size of concatenation is bounded by the sum of sizes $|P_{1}|+|P_{2}|+|T_{r,v_{1}}|+|C_{v_{1},v_{2}}|$ plus three additional non-terminals. ∎
5.4.6. The algorithm
Proposition 5.14.
Consider a segment $(P_{i-1},v_{i-1}),t^{-1},(P_{i},v_{i}),t,(P_{i+1},v_{i+1})$ in (7). There is an algorithm that in $O(|P_{i-1}|+|P_{i}|+|P_{i+1}|)$ time verifies if $w(P_{i},v_{i})\in H$ and, if so, constructs a syllable $(P,v)$ satisfying the following properties:
  • (a)
    $w(P,v)=_{G}w(P_{i-1},v_{i-1})\cdot t^{-1}w(P_{i},v_{i})t\cdot w(P_{i+1},v_{i+1})$,
  • (b)
    $|P|\leq|P_{i-1}|+|P_{i}|+|P_{i+1}|+C^{\ast}$.
The same holds for segments $(P_{i-1},v_{i-1}),t,(P_{i},v_{i}),t^{-1},(P_{i+1},v_{i+1})$.
Proof.
To verify if $w(P_{i},v_{i})\in H$ it is sufficient to check if $v_{i}$, viewed as an element of $F/N$, belongs to $H/N$. This can be precomputed for $\Gamma^{\ast}$, and hence can be checked in $O(1)$ time. Suppose that it is the case.
Using Proposition 5.12, one can compute in $O(|P_{i}|)$ time a syllable $(P^{\prime},v^{\prime})$ satisfying $w(P^{\prime},v^{\prime})=_{F}\varphi(w(P_{i},v_{i}))=_{G}t^{-1}w(P_{i},v_{i})t$. Then, using Proposition 5.13 twice, one can compute in $O(|P_{i-1}|+|P_{i}|+|P_{i+1}|)$ time concatenation $(P,v)$ of three syllables $(P_{i-1},v_{i-1})$, $(P^{\prime},v^{\prime})$, and $(P_{i+1},v_{i+1})$. Now, (a) follows from the construction of $(P,v)$. Furthermore,
$|P|\leq$ $|P_{i-1}|+|P^{\prime}|+|P_{i+1}|+2\bigg(\sum_{v}|T_{r,v}|+\sum_{v_{1},v_{2}}|C_{v_{1},v_{2}}|+3\bigg)$ (Proposition 5.13(b))
$\leq$ $|P_{i-1}|+|P_{i}|+|P_{i+1}|+2\bigg(\sum_{v}|T_{r,v}|+\sum_{v_{1},v_{2}}|C_{v_{1},v_{2}}|+3\bigg)$
$+\sum_{e\in E^{+}\setminus T}(|P_{e}^{\varphi}|+|P_{e^{-1}}^{\varphi}|)+\sum_{v}|P_{v}^{\varphi}|+1$ (Proposition 5.12(b))
$\leq$ $|P_{i-1}|+|P_{i}|+|P_{i+1}|+C^{\ast}$
and (b) holds. ∎
Theorem 5.15.
Suppose that $H_{1}=H_{2}$ are finite index subgroups of $F$ and let $\varphi:H_{1}\to H_{2}$ be normalizable. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.
Proof.
The proof is the same as the proof of Theorem 4.3, but instead of Proposition 4.2 we use Proposition 5.14 for a single Britton reduction step. ∎